OpenStudy (anonymous):

A circle has a radius of 6 inches. Find the area of a sector of this circle that is intercepted by a central angle measuring 30°. 2π 3π 6π 12π

OpenStudy (anonymous):

nice question again!! :) A central angle is measured by its intercepted arc. Let's denote the length of the intercepted arc with s, and the length of the radius r. So, s = 6 cm and r = 30 cm. When a central angle intercepts an arc whose length measure equals the length measure of the radius of the circle, this central angle has a measure 1 radian. To find the angle in our problem we use the following relationship: measure of an angle in radians = (length of the intercepted arc)/(length of the radius) measure of our angle = s/r = 6/30 = 1/5 radians. Now, we need to convert this measure angle in radians to degrees. Since pi radians = 180 degrees, then 1 radians = 180/pi degrees, so: 1/5 radians = (1/5)(180/pi) degrees = 36/pi degrees, or approximate to 11.5 degrees.

OpenStudy (anonymous):

so 6

OpenStudy (anonymous):

Thirty degrees is one twelfth of the disk, so find the area of the disk and divide by twelve. By mental math, I think the answer is about 3 pi.